Fermi National Laboratory

Accelerator Update

Friday September 24


    A0 Vacuum Upgrade - 60%

    D17 Separator Installation (discovered a vacuum leak) - 50%

    Conning Tower work at F2 & F3 - 50%

    E4 Vacuum Work - 80%

    IPM and Crystal Collimator installation - 20%

    D4 Spool Replacement - 80%

    Level Run - 100% complete (During a level run an alignment crew checks all the magnets positions relative to survey monuments.)

    Shimming 328 Dipoles - 52% complete (may add 84 more magnets to the list)

    Kaiser Coil Measurements - 90% complete (A Kaiser coil measurement checks to see if the cryo magnet's coil has twisted or shifted in any way.)

    Magnet Stand Replacement - Job canceled

    Unrolling Magnets - 20%


    Low Energy LCW System Upgrade - 100% complete

    Low Energy LCW Instrumentation Upgrade - 100% complete

    Power Amplifier Filament Power Supply LCW Hose Replacement - 100% complete

    High Energy RF Maintenance - 100% complete

    Mod VME Fans - 100% complete

    Drift Tube Tank #5 Replacement - 15% complete

    Drift Tube Tank #1 Vacuum Repairs - 100% complete

    Controls Maintenance - 80%


    Vacuum Work - 100% complete

    RF19 - 68% complete

    Long 13 Work - 78% complete (Long 13 is a sector of the Booster ring.)

    Vertical Collimator - 91% complete

Antiproton Source

    Debuncher Injection Area Beampipe and Quadrupole (Quad) replacement - The D20 & D30 pump down and leak check in progress.

      Debuncher Motorized Quads Installation - ~50% complete

      Debuncher Injection Region - The LBQs are set. The new injection kicker stand installation is progressing. They're waiting for beam pipe.

    Accumulator Flying Wires - The wires have been installed and tested. They are waiting on vacuum pump down.

    ARF1-1 Installation - ~90% complete. The leak check was successful. The final assembly is in progress.

    AP30 SO Cord Remediation - ~100% complete

    AP10 SO Cord Remediation - 50% complete

    Other Work - Target Station, Alignment, Stochastic Cooling, Controls, Cryo, LCW, LLRF - all work on track

Main Injector

      Q307-309 - Not started
      Q303-305 - ~100% complete

    MI-30 & 62


      Cooling Tank Repairs - ~55% complete


      Vacuum Modification - 100% complete

    Flying Wire Installation - ~100% complete

    Shielding Installation - 100% complete


    Horn Replacement - 52%

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 9/27/2004   email Fermilab